
While observing T0395113305 (00 47 30E -60 09 55) the telescope lost the track rate for a while. In order to re center the star in the fiber I turned off/on the track in the TCS, then moved back the star to the fiber. The star was out of the fiber for about 30 seconds.
Several times along the night I received from the GUI the message -- ASYNC ERROR: DHE warning: TSMTEMP2 bellow_allowed_range (0.000000)[temp]<<_chiron
I received four emails with the message -- Warning Timing chiron: table dhe_temps have not update in more than 120.0 minutes
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-03-04 | 10.0 | 0.0 | 0.0 | 0.0 | 0.0 | 10 |
Links
[1] mailto:thenry@astro.gsu.edu
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/36103/delete?destination=print/content/enr-ctio-15-m-2021-03-04
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/36103/edit?destination=print/content/enr-ctio-15-m-2021-03-04
[4] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/17569?destination=print/content/enr-ctio-15-m-2021-03-04
The program was completed successfully.